OpenStudy (anonymous):

Betsy has chosen a set of vases for her wedding registry. The vases are cylindrical and their dimensions (in inches) are as follows. • Vase A: h = 14 and r = 3 • Vase B: h = 7 and r = 6 Which of the following statements about the volumes of the vases is true? The volume of vase A is half the volume of vase B. The volume of vase B is half the volume of vase A. The volume of vase B is four times the volume of vase A. The volumes of the two vases are equal.

OpenStudy (anonymous):

@satellite73

OpenStudy (jdoe0001):

\(\large { \textit{volume of a cylinder}= \pi r^2 h \begin{cases} h=14\quad r=3\to \pi\cdot 14^2\cdot 3 \\ \quad \\ h=7\quad r=6\to \pi \cdot 7^2\cdot 6 \end{cases} \\ \quad \\ \cfrac{\pi\cdot 14^2\cdot 3}{\pi \cdot 7^2\cdot 6}\implies \cfrac{\pi\cdot (2\cdot 7)^2\cdot 3}{\pi \cdot 7^2\cdot 2\cdot 3} \implies \cfrac{\pi\cdot (2\cdot 7)(2\cdot 7)\cdot 3}{\pi \cdot 7\cdot 7\cdot2\cdot 3} \\ \quad \\ \cfrac{\pi\cdot 2\cdot 7\cdot 2\cdot 7\cdot 3}{\pi \cdot 7\cdot 7\cdot2\cdot 3} }\) what do you think?

OpenStudy (jdoe0001):

volume of one over the other would give you the ratio of "one" over the "other"

OpenStudy (anonymous):

c or b i think

OpenStudy (jdoe0001):

well, what does the fraction simplify to?

OpenStudy (anonymous):

c i think

OpenStudy (jdoe0001):

well, the fraction has no variable "c", so... it wouldn't simplify to that

OpenStudy (anonymous):

Im confused

jimthompson5910 (jim_thompson5910):

what is the volume of vase A? use the formula V = pi*r^2*h

OpenStudy (jdoe0001):

you're asked on what is the "ratio" of one of the vases in relation to the other is it half, one third whatever else thus is a ratio matter, or, what is the ratio of one vase in relation to the other thus "a" in relation to "b" would be a:b or \(\frac{a}{b}\)

OpenStudy (anonymous):

okay